# On irreducibility of six-dimensional compatible systems of $\mathbb{Q}$

### v1

@@ 3.1.3. $[K:\mathbb{Q}]=2$
Suppose $\rho_{\lambda}$ is reducible for infinitely many $\lambda$. By 2.24, one can assume each such $\rho_{\lambda}$ contains a dimension $\geq 3$ irreducible component $W_{\lambda}$. Suppose first $\dim W_{\lambda}=3$, then the irreducible decomposition of $\rho_{\lambda}$ has to be
$\rho_{\lambda}=W_{\lambda}\oplus W_{\lambda}^{\prime}$
with $\dim W_{\lambda}^{\prime}=3$ since otherwise would contradict with 2.15. If $W_{\lambda}\otimes\chi\cong W_{\lambda}^{\prime}$, since $\chi$ has finite image, after restricting to a finite extension, the two components $W_{\lambda}$ and $W_{\lambda}^{\prime}$ are isomorphic contradicts with regularity of the compatible system. Hence one has $W_{\lambda}\cong W_{\lambda}\otimes\chi$. Twist $\rho_{\lambda}$ with a suitable power of cyclotomic character, one can assume $W_{\lambda}$ has algebraic monodromy group $\text{G}_{W_{\lambda}}=\text{GO}_{3}$ or $\text{GL}_{3}$. But this would contradict with $W_{\lambda}\cong W_{\lambda}\otimes\chi$.
@@
Assume $\dim W_{\lambda}=4$ or 5. By 2.16, one can assume each of the irreducible components of $\rho_{\lambda}$ cannot be Lie-irreducible. We also assume not every irreducible component of $\rho_{\lambda}$ is induced by some character since otherwise we use method in 3.1.1 to contradict with $\rho_{\lambda_{0}}$ being irreducible. Then the only remaining case is:
$\rho_{\lambda}=\text{Ind}^{\mathbb{Q}}_{F}\sigma_{\lambda}\oplus\text{Ind}^{\mathbb{Q}}_{F}\chi_{\lambda}$
@@
where $\sigma_{\lambda}$ is 2-dimensional Lie-irreducible of $F$, and $\chi_{\lambda}$ is a character of $F$, and $F$ is a quadratic number field. For $\lambda$ sufficiently large, the 2-dimensional part $W_{\lambda}^{\prime}$ fits into a strictly compatible system by 2.23.(i). We will show the four-dimensional part is potential automorphic. Then the semisimplicity of compatible system would contradict the irreducibility of $\rho_{\lambda_{0}}$.
It is enough to check the conditions in 2.21. (i) is obvious, (iii) can be obtained from [BLGGT14, Lemma 1.4.3.(2)]. To check (ii), as $\sigma_{\lambda}$ is Lie-irreducible, the algebraic monodromy group of $W_{\lambda}$ must contain $\text{SL}_{2}\times\text{SL}_{2}$. Since $\rho_{\lambda}$ satisfies $\rho_{\lambda}\cong\rho_{\lambda}^{\vee}\otimes\chi_{\lambda}$, by dimensional reason one has $W_{\lambda}\cong W_{\lambda}^{\vee}\otimes\chi_{\lambda}$. Hence the algebraic monodromy group of $W_{\lambda}$ must factor through $\text{GSp}_{4}$. Moreover, one can always twist $\chi$ to $\{\chi_{\lambda}\}$ making them odd. Hence $W_{\lambda}$ is essentially self-dual and totally odd. Finally, we check (iv), namely for $\lambda$ suffciently large, $\tau=(\text{Ind}_{F}^{\mathbb{Q}}\sigma_{\lambda})|_{\mathbb{Q}(\zeta_{\ell})}$ is residually irreducible, where $\ell$ is the unique rational prime divided by $\lambda$.
For $\ell$ sufficiently large, $\mathbb{Q}(\zeta_{\ell})$ and $F$ are linearly disjoint. Hence one has:
$\text{Res}_{\mathbb{Q}(\zeta_{\ell})}\text{Ind}_{F}^{\mathbb{Q}}\sigma_{\lambda}=\text{Ind}_{F(\zeta_{\ell})}^{\mathbb{Q}(\zeta_{\ell})}\text{Res}_{F(\zeta_{\ell})}\sigma_{\lambda}$
@@
As $\sigma_{\lambda}$ is Lie-irreducible, $\text{Res}_{F(\zeta_{\ell})}\sigma_{\lambda}$ is irreducible. But this restriction is of type A since it has dimension 3, hence the induced representation also is of type A. Hence by 2.11.(iv) it is residually irreducible.
@@
In this non Lie-irreducible case, our proof also works for compatible systems over totally real field, except that to apply potential automorphy theorem in 3.1.3, one has to assume for all complex conjugation $c$, the similitude character $\chi_{\lambda_{0}}(c)$ has same sign. Note that 1 also is true for totally real field, and under above extral condition 2 also is true for totally real fields. We summarize the result as following.
Theorem 3.3.
Let $\{\rho_{\lambda}:\mathrm{Gal}_{\mathbb{Q}}\to\text{GL}_{6}(\overline{E_{\lambda}})\}$ be a 6-dimensional $E$-rational regular strictly compatible system of $\mathbb{Q}$. Suppose some $\rho_{\lambda_{0}}$ is irreducible but not Lie-irreducible. Suppose further that $\rho_{\lambda_{0}}\cong\rho_{\lambda_{0}}^{\vee}\otimes\chi$ and the sign $\chi(c)$ is independent of complex conjugation $c$, then $\rho_{\lambda}$ is irreducible for all but finitely many $\lambda$.
@@ 3.2. The Lie-irreducible case
Here symbol $\text{Sym}^{r}(\sigma)$ denotes the $r$th symmetric power of a representation $\sigma$, and Std stands for the standard representation. We prove the following result:
@@
    Assume $\rho_{\lambda_{0}}$ has semisimple type (b), then each $\rho_{\lambda}$ is Lie-irreducible with same semisimple type for all but finitely many $\lambda$.
  3. (iii)
    Assume $\rho_{\lambda_{0}}$ is fully polarisable (this means (c) or (d)), then each $\rho_{\lambda}$ is Lie-irreducible with same semisimple type for all but finitely many $\lambda$.
@@ 3.2.3. Cases (c) and (d)
Next we distinguish the semisimple type $\text{Sp}_{6}$ and $\text{SO}_{6}$. Assume our compatible system $\{\rho_{\lambda}\}$ have infinitely many terms with semisimple type $\text{SO}_{6}$ and infinitely many terms with semisimple type $\text{Sp}_{6}$. If the similitude character $\chi_{\lambda_{0}}$ is odd (resp. even), one can choose all similitude characters $\chi_{\lambda}$ of $\rho_{\lambda}$ to be odd (resp. even). Assume $\rho_{\lambda}$ belongs to either one of the following infinite sets:
@@
We only need to check conditions in 2.21. And it is only condition (iv) requires further explaination. We first show $\overline{\rho_{\lambda}}^{\text{ss}}$ is irreducible for all but finitely many $\lambda$. This also is partial proof of 1.3.(iii). Note if the semisimple type is $\text{SO}_{6}$ which is type A, by 2.11.(iii) $\overline{\rho}_{\lambda}^{\text{ss}}$ is irreducible when restricting to $\mathbb{Q}(\zeta_{\ell})$ for all but finitely many $\lambda$.
So we are in case (i) above. As the formal bi-character of $\rho_{\lambda}$ and $\overline{\rho}_{\lambda}^{\text{ss}}$ coincide. If the latter is reducible, there exists a 2-dimensional irreducible component $\overline{\sigma}_{\lambda}$. Oddness and regularity guaratee this $\sigma_{\lambda}$ is odd. Hence by 2.25.(i), there exists a 2-dimensional strictly compatible system $\{\varphi_{\lambda}\}$ such that for infinitely many $\lambda$ one has $\overline{\varphi}_{\lambda}^{\text{ss}}\cong\overline{\sigma}_{\lambda}$. Construct an 8-dimensional strictly compatible system:
@@
Then on the one hand by Goursat’s lemma $\textbf{G}_{\lambda}^{\text{der}}=\text{Sp}_{6}\times\text{SL}_{2}$ which has rank 4. On the other hand for some $\lambda$ one has $\overline{\psi}_{\lambda}^{\text{ss}}=\overline{\rho}_{\lambda}^{\text{ss}}\oplus\overline{\sigma}_{\lambda}$ which by 2.7 has rank 3. This is a contradiction.
Next we show $\overline{\rho}_{\lambda}^{\text{ss}}|_{\text{Gal}_{\mathbb{Q}(\zeta_{\ell})}}$ is irreducible. If on the contrary this is reducible, by [CG13, Lemma 4.3] $\overline{\rho}_{\lambda}^{\text{ss}}$ is of induced type. Hence the Dirichlet density $d(\overline{\rho}_{\lambda}^{\text{ss}})$ is strictly smaller than 1. However, by [Pa19, Proposition 3.4.9.(i)] this density is the Dirichlet density of the set of primes whose Frobenius has zero trace. But since $\rho_{\lambda}$ is Lie-irreducible, this density should be 1, a contradiction.
@@ 3.2.4. Case (b)
Since the formal bi-character is $\{x,x^{-1},xy,x^{-1}y,xy^{-1},x^{-1}y^{-1}\}$. If $\rho_{\lambda}$ is reducible, the irreducible decomposition has to be $\rho_{\lambda}=W_{\lambda}\oplus W_{\lambda}^{\prime}$ with $\dim W_{\lambda}=2,\dim W_{\lambda}^{\prime}=4$. As $\text{Sp}_{4}$ does not have 2-dimensional irreducible representations, $\textbf{G}_{W_{\lambda}^{\prime}}^{\text{der}}$ has to be $\text{SO}_{4}$. Hence the representation factors through $\text{GO}_{4}$.
@@
We show for $\lambda$ sufficiently large, both $\sigma_{\lambda}$ and $\sigma_{\lambda}^{\prime}$ fit into compatible systems. Then by semisimplicity this contradicts with $\rho_{\lambda_{0}}$ being irreducible.
It is enough to assume both $\sigma_{\lambda}$ and $\sigma_{\lambda}^{\prime}$ are Lie-irreducible. We check the conditions in 2.21. (i) is obvious, (iii) as before follows from [BLGGT14, Lemma 1.4.3.(2)]. (iv) follows from 2.11.(iii). We use 2.22 to check the oddness condition (ii). In those four conditions, (i) and (iii) are obvious. (ii) follows from 2.11.(iii). Finally for the boundness of difference of Hodge-Tate weights, this comes from the facts that Hodge-Tate weights of all $\rho_{\lambda}$ are the same, and the set of differences of Hodge-Tate weights of $\sigma_{\lambda}$ and $\sigma_{\lambda}^{\prime}$ are contained in the set of differences of Hodge-Tate weights of $\rho_{\lambda}$.
@@
In our proof of 1.3(i), everything works for a regular strictly compatible system of a general totally real field $F$, except for this case (b). Here we need to write a Galois representation $\text{Gal}_{F}\to\text{GO}_{4}(\overline{E_{\lambda}})$ as the tensor product of two 2-dimensional representations with certain $\ell$-adic hodge theoretical properties, which can be done for $F=\mathbb{Q}$. If however the strictly compatible system of a totally real field is pure, then by [Pa19, Section 3.2], one still have (1) with the desired properties. Then our proof works fine for such compatible systems.
3.3. Residual irreducibility
We summarize the proof of 1.3.(iii). The non Lie-irreducible cases can be proven the same as in the end of 3.1.3. For the Lie-irreducible cases, type A ones follow from 2.11.(iii); the fully symplectic case can be proven the same as in 3.2.3.
@@ 3.3. Residual irreducibility
As we are assuming purity of the compatible system, 4 asserts 1.3(i) is true for compatible systems of totally real fields. Hence we can assume each $\rho_{\lambda}$ is reducible. We first prove two lemmas:
@@
For all but finitely many $\lambda$, the irreducible component of $\rho_{\lambda}$ which either has dimension 1,2,4 or has dimension 3 with semisimple type $\text{SO}_{3}$, fits into a strictly compatible system.
@@
Assume $\dim\sigma_{\lambda}=4$. This is essentially self-dual and totally odd due to dimensional resason. Hence to apply 2.21 we only need to check $\overline{\sigma}_{\lambda}^{\text{ss}}|_{\text{Gal}_{\mathbb{Q}(\zeta_{\ell})}}$ is irreducible. If it is induced, then the same method as in 3.1.3 shows $\overline{\sigma}_{\lambda}^{\text{ss}}|_{\text{Gal}_{\mathbb{Q}(\zeta_{\ell})}}$ is irreducible. If it is Lie-irreducible, if the semisimple type is of type A, then 2.11.(iii) shows the desired result. Hence we only need to assume the semisimple type of $\sigma_{\lambda}$ is fully symplectic. Replace $\text{Sp}_{6}$ with $\text{Sp}_{4}$, the same proof as in 3.2.3 applies. ∎
@@
Suppose on the contrary such $W_{\lambda_{0}}$ exists. By essential self-duality, one must have irreducible decomposition $\rho_{\lambda_{0}}=W_{\lambda_{0}}\oplus W^{\prime}_{\lambda_{0}}$ with $W^{\prime}_{\lambda_{0}}=W^{\vee}_{\lambda_{0}}\otimes\chi_{\lambda_{0}}$. Hence there cannot be infinitely many $\lambda$ such that the irreducible components of $\rho_{\lambda}$ are all of types mentioned in 3.5. (If there are infinitely many $\lambda$ such that $\rho_{\lambda}$ has dimensional type 3+3 and each component has semisimple type $\text{SO}_{3}$, then those components extends to a compatible system and contains $W_{\lambda_{0}}$. The semisimple ranks do not match).
Hence the only possible dimensional type of $\rho_{\lambda}$ other than 3+3 is 5+1. Denote by $d$ the Dirichlet density of $\lambda$ such that $\rho_{\lambda}$ has dimensional type 5+1. If $d>0$, then by 2.26 there are infinitely many those 5-dimensional component satisfies 2.21.(iv). Other conditions in 2.21 are easy to verify. Hence one writes $\{\rho_{\lambda}\}$ as a direct sum of a 5-dimensional compatible system and a 1-dimensional compatible system. This contradicts with the decomposition of $\rho_{\lambda_{0}}$. So we must have $d=0$. But then the Dirichlet density of $\lambda$ such that $\rho_{\lambda}$ contains a 3-dimensional component with semisimple type $\text{SL}_{3}$ would be 1. This contradicts with [PT15, Lemma 1.6]. ∎
To finish the proof, denote by $r$ the smallest integer such that there are infinitely many $\lambda$ such that each the irreducible components of $\rho_{\lambda}$ have dimenion $\leq r$. Then 3.5 asserts the result is true for $r\leq 4$. Assume $r=5$, then one again uses 2.26 and checks conditions in 2.21 to show the 5-dimensional components extend to a strictly compatible system. This finishes the proof.

### v3

@@ 3.1.3. $[K:\mathbb{Q}]=2$
Suppose $\rho_{\lambda}$ is reducible for infinitely many $\lambda$. By 2.16 each conponent of $\rho_{\lambda}$ is not Lie-irreducible. In particular, there cannot be one-dimensional components. We also assume not every irreducible component of $\rho_{\lambda}$ is induced by some character since otherwise we use method in 3.1.1 to contradict with $\rho_{\lambda_{0}}$ being irreducible. The only remaining case is:
$\rho_{\lambda}=\text{Ind}^{\mathbb{Q}}_{F_{1}}\sigma_{\lambda}\oplus\text{Ind}^{\mathbb{Q}}_{F_{2}}\chi_{\lambda}$
where $\sigma_{\lambda}$ is 2-dimensional Lie-irreducible of $F_{1}$, and $\chi_{\lambda}$ is a character of $F_{2}$, and $F_{1},F_{2}$ are quadratic number fields. As $d(\rho_{\lambda_{0}})=1/2$, by 2.15 and 3.2 we have $F_{1}=F_{2}$. And in fact one uses method in 1 to show this quadratic field is $K$. Moreover $K$ is an imaginary quadratic field otherwise would contract regularity of $\text{Ind}_{K}^{\mathbb{Q}}\chi_{\lambda}$.
@@
By class field theory, the 2-dimensional part $W_{\lambda}^{\prime}$ always fits into a strictly compatible system for $\lambda$ sufficiently large. We will show the four-dimensional part $W_{\lambda}$ is potential automorphic for $\lambda$ sufficiently large. Then the semisimplicity of compatible system would contradict the irreducibility of $\rho_{\lambda_{0}}$.
It is enough to check the conditions in 2.21. (i) is obvious, (iii) can be obtained from [BLGGT14, Lemma 1.4.3.(2)]. To check (ii), we first show $W_{\lambda}$ factors through $\text{GSp}_{4}$. To do so it is enough to show the actions of $\text{Gal}_{F}$ and a fixed complex conjugation $c$ of $\mathbb{Q}$ factor through $\text{GSp}_{4}$. Restrict $W_{\lambda}$ to $K$ gives a direct sum of two 2-dimensional Lie-irreducible representations of $F$. Compare to the first component, the action on the second one is conjugate by $c$. Hence this restriction factors through $\mathbb{G}_{m}\left(\text{SL}_{2}\times\text{SL}_{2}\right)$. This factors through $\text{GSp}_{4}$ since the standard action of $\text{SL}_{2}\times\text{SL}_{2}$ on $\overline{E_{\lambda}}^{2}\oplus\overline{E_{\lambda}}^{2}$ preserve the direct sum of the two determinant forms. The complex conjugation swaps the two irreducible components and also preserve this alternating form, hence is also contained in $\text{GSp}_{4}$. To deal with oddness, as $W_{\lambda}\cong W_{\lambda}\otimes\chi$ where $\chi$ is the character at the beginning, hence one can always make similitude character odd after possibly twisting $\chi$. Hence $W_{\lambda}$ is essentially self-dual and totally odd.
Finally, we check (iv), namely for $\lambda$ suffciently large, $\tau=(\text{Ind}_{K}^{\mathbb{Q}}\sigma_{\lambda})|_{\mathbb{Q}(\zeta_{\ell})}$ is residually irreducible, where $\ell$ is the unique rational prime divided by $\lambda$. For $\ell$ sufficiently large, $\mathbb{Q}(\zeta_{\ell})$ and $K$ are linearly disjoint. Hence one has:
$\text{Res}_{\mathbb{Q}(\zeta_{\ell})}\text{Ind}_{K}^{\mathbb{Q}}\sigma_{\lambda}=\text{Ind}_{K(\zeta_{\ell})}^{\mathbb{Q}(\zeta_{\ell})}\text{Res}_{K(\zeta_{\ell})}\sigma_{\lambda}$
@@
As $\sigma_{\lambda}$ is Lie-irreducible, $\text{Res}_{K(\zeta_{\ell})}\sigma_{\lambda}$ is irreducible. But this restriction is of type A since it has dimension 3, hence the induced representation also is of type A. Hence by 2.11.(iv) it is residually irreducible.
@@
In this non Lie-irreducible case, our proof also works for compatible systems over totally real field, except that to apply potential automorphy theorem in 3.1.3, one has to assume for all complex conjugation $c$, the similitude character $\chi_{\lambda_{0}}(c)$ has same sign. Note that 1 also is true for totally real fields, and under above extral condition, 2 also is true for totally real fields.
@@ 3.2. The Lie-irreducible case
Here symbol $\text{Sym}^{r}(\sigma)$ denotes the $r$th symmetric power of a representation $\sigma$, and Std denotes the standard representation. We prove the following result:
@@
    Assume $\rho_{\lambda_{0}}$ has semisimple type (b) or is fully polarisable (this means (c) or (d)), then $\rho_{\lambda}$ is Lie-irreducible with same semisimple type for all but finitely many $\lambda$.
@@ 3.2.3. Cases (c) and (d)
Next we distinguish the semisimple type $\text{Sp}_{6}$ and $\text{SO}_{6}$. Assume our compatible system $\{\rho_{\lambda}\}$ have infinitely many terms with semisimple type $\text{SO}_{6}$ and infinitely many terms with semisimple type $\text{Sp}_{6}$. If the similitude character $\chi_{\lambda_{0}}$ is odd (resp. even), by 2.19 one can choose all similitude characters $\chi_{\lambda}$ of $\rho_{\lambda}$ to be odd (resp. even). Assume $\rho_{\lambda}$ belongs to either one of the following infinite sets:
@@
We only need to check conditions in 2.21. And it is only condition (iv) requires further explaination. We first show $\overline{\rho_{\lambda}}^{\text{ss}}$ is irreducible for all but finitely many $\lambda$ in the infinite set above. Note if the semisimple type is $\text{SO}_{6}$ which is of type A, by 2.11.(iii) $\overline{\rho}_{\lambda}^{\text{ss}}$ is irreducible when restricting to $\mathbb{Q}(\zeta_{\ell})$ for all but finitely many $\lambda$.
So we are in case (i) above. As the formal bi-character of $\rho_{\lambda}$ and $\overline{\rho}_{\lambda}^{\text{ss}}$ coincide. If the latter is reducible, there exists a 2-dimensional irreducible component $\overline{\sigma}_{\lambda}$. Oddness and regularity of $\rho_{\lambda}$ guaratee this $\sigma_{\lambda}$ is odd. Hence by 2.25.(i), after possibly twisting some power of cyclotomic character to the whole compatible system, there exists a 2-dimensional strictly compatible system $\{\varphi_{\lambda}\}$ such that for infinitely many $\lambda$ one has $\overline{\varphi}_{\lambda}^{\text{ss}}\cong\overline{\sigma}_{\lambda}$. Construct an 8-dimensional strictly compatible system:
@@
Then on the one hand by Goursat’s lemma $\textbf{G}_{\lambda}^{\text{der}}=\text{Sp}_{6}\times\text{SL}_{2}$ which has rank 4. On the other hand for some $\lambda$ one has $\overline{\psi}_{\lambda}^{\text{ss}}=\overline{\rho}_{\lambda}^{\text{ss}}\oplus\overline{\sigma}_{\lambda}$. The algebraic envelope has semisimple rank 3, which contradicts with 2.11.(i).
Next we show $\overline{\rho}_{\lambda}^{\text{ss}}|_{\text{Gal}_{\mathbb{Q}(\zeta_{\ell})}}$ is irreducible for all but finitely many $\lambda$ in the infinitely set (i) above. If on the contrary this is always reducible, by [CG13, Lemma 4.3] $\overline{\rho}_{\lambda}^{\text{ss}}$ is induced by a 2-dimensional Lie-irreducible representation of some quadratic field $K_{\lambda}$ which is contained in $L$ defined by 2.11.(ii). Hence there are infinitely many such $\lambda$ with $K_{\lambda}$ coincide, denote by $K$. Then the trace of rational primes that the corresponding Frobenius element is the nontrivial element of $\text{Gal}(K/\mathbb{Q})$ is zero. This provides that the Dirichlet density of trace zero primes under $\rho_{\lambda}$ is at least 1/2. However, since $\rho_{\lambda_{0}}$ is Lie-irreducible, the density should be zero, a contradiction.
@@ 3.2.4. Case (b)
Suppose there are infinitely many $\lambda$ such that $\rho_{\lambda}$ is reducible. Since the formal bi-character is $\{x,x^{-1},xy,x^{-1}y,xy^{-1},x^{-1}y^{-1}\}$. If $\rho_{\lambda}$ is reducible, the irreducible decomposition has to be $\rho_{\lambda}=W_{\lambda}\oplus W_{\lambda}^{\prime}$ with $\dim W_{\lambda}=2,\dim W_{\lambda}^{\prime}=4$. The 2-dimensional part $W_{\lambda}$ fits into a strictly compatible system for all but finitely many $\lambda$. As $\text{Sp}_{4}$ does not have 2-dimensional irreducible representations, $\textbf{G}_{W_{\lambda}^{\prime}}^{\text{der}}$ has to be $\text{SO}_{4}$. Hence the representation factors through $\text{GO}_{4}$.
@@
We show for $\lambda$ sufficiently large, both $\sigma_{\lambda}$ and $\sigma_{\lambda}^{\prime}$ fit into compatible systems. Then both $W_{\lambda}$ and $W_{\lambda}^{\prime}$ fit into strictly compatible systems for $\lambda$ sufficiently large. Then by semisimplicity this contradicts with $\rho_{\lambda_{0}}$ being irreducible.
It is enough to assume both $\sigma_{\lambda}$ and $\sigma_{\lambda}^{\prime}$ are Lie-irreducible. We check the conditions in 2.21. (i) is obvious, (iii) as before follows from [BLGGT14, Lemma 1.4.3.(2)]. (iv) follows from 2.11.(iii). We use 2.22 to check the oddness condition (ii). In those four conditions, (i) and (iii) are obvious. (ii) follows from 2.11.(iii). Finally for the boundness of difference of Hodge-Tate weights, this comes from the facts that $\tau$-Hodge-Tate weights of all $\rho_{\lambda}$ are the same, and the set of differences of Hodge-Tate weights of $\sigma_{\lambda}$ and $\sigma_{\lambda}^{\prime}$ are contained in the set of differences of Hodge-Tate weights of $\rho_{\lambda}$.
@@
In our proof of 1.3(i), everything works for a regular strictly compatible system of a general totally real field $F$, except for this case (b). Here we need to write a Galois representation $\text{Gal}_{F}\to\text{GO}_{4}(\overline{E_{\lambda}})$ as the tensor product of two 2-dimensional representations with certain $\ell$-adic hodge theoretical properties, which can be done for $F=\mathbb{Q}$. If however the strictly compatible system of a totally real field is pure, then by [Pa19, Section 3.2] (see also [DWW24, Section 2.3] and [PT15, Lemma 1.2]), one still has (2) with the desired properties. Then our proof works fine for such compatible systems. We summarize the result for totally real fields in below. Note also that in order to make our proof works for 3.3.(ii) of totally real fields, more precisely the statement that fully polarisable compatible systems have same semisimple type for all but finitely many representations, one has to add a condition that the similitude character of $\rho_{\lambda_{0}}$ is totally odd or totally even.
Theorem 3.4.
Let $\{\rho_{\lambda}:\mathrm{Gal}_{K}\to\text{GL}_{6}(\overline{E_{\lambda}})\}$ be a 6-dimensional $E$-rational pure regular strictly compatible system of a totally real field $K$. Suppose some $\rho_{\lambda_{0}}$ is Lie-irreducible. Then $\rho_{\lambda}$ is (Lie-)irreducible for all but finitely many $\lambda$.
@@
By 1.3 we assume each $\rho_{\lambda}$ is reducible. We first prove two lemmas:
@@
For all but finitely many $\lambda$, the irreducible component of $\rho_{\lambda}$ which either has dimension 1,2 or has dimension 3 with semisimple type $\text{SO}_{3}$, fits into a strictly compatible system. If there are infinitely many $\lambda$ such that $\rho_{\lambda}$ has a 4-dimensional irreducible component, then some of them fits into a strictly compatible system.
@@
Assume $\dim\sigma_{\lambda}=4$. This is essentially self-dual and totally odd due to dimensional resason. Hence to apply 2.21 we only need to check $\overline{\sigma}_{\lambda}^{\text{ss}}|_{\text{Gal}_{\mathbb{Q}(\zeta_{\ell})}}$ is irreducible. If $\sigma_{\lambda}$ is induced, then the same method as in 3.1.3 shows $\overline{\sigma}_{\lambda}^{\text{ss}}|_{\text{Gal}_{\mathbb{Q}(\zeta_{\ell})}}$ is irreducible. If it is Lie-irreducible, if the semisimple type is of type A, then 2.11.(iii) shows the desired result. Hence we only need to assume the semisimple type of $\sigma_{\lambda}$ is fully symplectic. Replace $\text{Sp}_{6}$ with $\text{Sp}_{4}$, the same proof as in 3.2.3 applies. ∎
@@
Suppose on the contrary such $W_{\lambda_{0}}$ exists. By essential self-duality, one must have irreducible decomposition $\rho_{\lambda_{0}}=W_{\lambda_{0}}\oplus W^{\prime}_{\lambda_{0}}$ with $W^{\prime}_{\lambda_{0}}=W^{\vee}_{\lambda_{0}}\otimes\chi_{\lambda_{0}}$. Hence there cannot be infinitely many $\lambda$ such that the irreducible components of $\rho_{\lambda}$ are all of types mentioned in 3.5. (If there are infinitely many $\lambda$ such that $\rho_{\lambda}$ has dimensional type $3+3$ and each component has semisimple type $\text{SO}_{3}$, then those components extends to a compatible system which contains $W_{\lambda_{0}}$. The semisimple ranks do not match).
Hence for all but finitely many $\lambda$, the only possible dimensional type of $\rho_{\lambda}$ other than $3+3$ is $5+1$. Denote by $d$ the Dirichlet density of $\lambda$ such that $\rho_{\lambda}$ has dimensional type $5+1$. If $d>0$, then by 2.26 there are infinitely many those 5-dimensional component satisfies 2.21.(iv). Other conditions in 2.21 are easy to verify. Hence one writes $\{\rho_{\lambda}\}$ as a direct sum of a 5-dimensional compatible system and a 1-dimensional compatible system. This contradicts with the decomposition of $\rho_{\lambda_{0}}$. So we must have $d=0$. But then the Dirichlet density of $\lambda$ such that $\rho_{\lambda}$ contains a 3-dimensional component with semisimple type $\text{SL}_{3}$ would be 1. This contradicts with [PT15, Lemma 1.6]. ∎
To finish the proof, denote by $r$ the smallest integer such that there are infinitely many $\lambda$ such that each irreducible component of $\rho_{\lambda}$ has dimension $\leq r$. Then 3.5 and 3.6 assert the result is true for $r\leq 4$. Assume $r=5$, then one again checks conditions in 2.21 (in particular again use 2.26 to check (iv).) to show the 5-dimensional components extend to a strictly compatible system. This finishes the proof.